Comerpickles:

The radius of a cylinder is increased by a factor of 5. The height is not changed. What is the effect of the increase on the volume of the cylinder? A) The volume of the cylinder increases by a factor of 5. Eliminate B) The volume of the cylinder increases by a factor of 25. C) The volume of the cylinder increases by a factor of 15. D) The volume of the cylinder increases by a factor of 10.

Comerpickles:

@tigerlover

tigerlover:

We start with the surface area equation, S = 2(pi)r2 + 2(pi)rh, if the height and radius were doubled then the knew equation becomes S' = 2(pi)(2r)2 + 2(pi)(2r)(2h) = 4(2(pi)r2 + 4(2(pi)rh) = 4(2(pi)r2 +2(pi)rh). Examining the final result carefully we notice that the knew surface area S' = 4S so it increases by 4 times its orginal.

Comerpickles:

So what do I do next, I mean how do I determine the volume of the cylinder of whatever factor it increases by

tigerlover:

"Examining the final result carefully we notice that the knew surface area S' = 4S so it increases by 4 times its orginal."

Comerpickles:

So the answer would be A?

tigerlover:

Yes

Comerpickles:

Okie, can you help me with 8 more please?